# MediSwift: Efficient Sparse Pre-trained Biomedical Language Models

### v1

@@ 4 Fine-tuning on Biomedical tasks
This section evaluates MediSwift’s performance on PubMedQA and HoC benchmarks using dense fine-tuning on each variant’s specific training set (e.g., fine-tuning and testing with the PubMedQA dataset). During the fine-tuning phase, the total training FLOPs constitute only a minor fraction of the FLOPs expended during pre-training, despite the fine-tuning being conducted densely. As a result, the FLOPs consumed during sparse pre-training is proportional to the combined FLOPs of sparse pre-training and dense fine-tuning.
Following Luo et al. (2022), we incorporate soft prompting into our fine-tuning framework by formatting sequences as [source; prompt; target]. This format helps our models better utilize contextual information, demonstrating MediSwift’s effectiveness and adaptability in medical text analysis. Further details on hyperparameters and dataset specifics are provided in Appendix A.2 and A.3. We note that all fine-tuning results were averaged across 3 random seeds.
@@ 4.2 Document Classification on HoC
The performance is evaluated using the micro-F1 score, allowing direct comparison with prior models and demonstrating our method’s effectiveness. In Table 4, the dense MediSwift-XL model outperformed all similarly sized models in micro-F1 score, with its 50% and 75% sparse demonstrating very competitive results, emphasizing sparse pre-training’s balance of computational efficiency and accuracy. This further showcases the potential of sparsity in optimizing language model performance for biomedical applications.
@@ 5 Related Work
Sparse weight training for language models has emerged as a promising avenue to address the computational intensity of training large models. Recent studies have explored various sparse training methodologies (Thangarasa et al., 2023a; Dao et al., 2022a; Chen et al., 2022), aiming to maintain or enhance model performance while significantly reducing computational requirements. Techniques such as pruning (Chen et al., 2020), sparse activations (Mirzadeh et al., 2024), along with the development of specialized software (NeuralMagic, 2021; Gupta, 2024) and hardware (Lie, 2022b, a; Dietrich et al., 2021) have been pivotal. Our work builds on these foundations, focusing on optimizing sparse weight training strategies specifically for domain-specific language models, pushing the boundaries of efficiency.
@@
The evolution of language models for medical applications has progressed from adapting encoder-only architectures like BERT (Devlin et al., 2018), using biomedical corpora (Lee et al., 2019; Gu et al., 2021), to incorporating strategies like document links (Yasunaga et al., 2022b) and knowledge graphs (Yasunaga et al., 2022a). The shift towards autoregressive generative models, such as GPT (Brown et al., 2020b) and Llama (Touvron et al., 2023a), for pretraining on medical texts has led to significant advancements (Wu et al., 2023a; Luo et al., 2022; Bolton et al., 2021). Recent scaling efforts include GatorTronGPT with 20B parameters (Yang et al., 2022), as well as Clinical-Camel (Toma et al., 2023), MEDITRON (Chen et al., 2023) and Med-42 (Christophe et al., 2023), based on Llama-2-70B (Touvron et al., 2023b), focusing on mixed clinical and general English texts.
Our work diverges from works that scaled up medical LMs by introducing weight sparsity into the pre-training of biomedical LMs. This reduces the computational costs typically associated with large-scale models, thereby improving the balance between efficiency and accuracy in the medical domain.
@@ 6 Conclusion
7 Limitations
Our work on MediSwift represents a significant leap forward in developing efficient domain-specific LLMs, particularly in biomedicine, by utilizing sparse pre-training to strike a fine balance between computational efficiency and accuracy. While we have initially focused on static sparse pre-training, the emerging field of dynamic sparse training (DST) holds great promise for further improvements (Evci et al., 2020; Mocanu et al., 2018; Liu et al., 2021).

### v2

@@ 4 Fine-tuning on Biomedical tasks
This section evaluates MediSwift’s performance on PubMedQA and HoC benchmarks using dense fine-tuning on each variant’s specific training set (e.g., fine-tuning and testing with the PubMedQA dataset). During the fine-tuning phase, the total training FLOPs constitute only a minor fraction of the FLOPs expended during pre-training, despite the fine-tuning being conducted densely. As a result, the FLOPs consumed during sparse pre-training is proportional to the combined FLOPs of sparse pre-training and dense fine-tuning. Following Luo et al. (2022), we incorporate soft prompting into our fine-tuning framework by formatting sequences as [source; prompt; target]. This format helps our models better utilize contextual information, demonstrating MediSwift’s effectiveness and adaptability in medical text analysis. Further details on hyperparameters and dataset specifics are provided in Appendix A.2 and A.3. We note that all fine-tuning results were averaged across 3 random seeds.
@@ 4.1 Question Answering with PubmedQA
Prompt Tuning Ablation Study
To assess the effectiveness of soft prompting, we have conducted the requested analysis and provide the results below. Our study focused on evaluating the impact of soft prompting on the performance of our MediSwift models during the fine-tuning phase, across both dense and sparse configurations (see Table 1). The approach involves steering the pre-trained language model by appending several additional virtual tokens as prompts before the text. These continuous embeddings, distinct from the main text, are randomly initialized and learned end-to-end on downstream tasks, making them task-specific. Unlike prefix tuning (Li and Liang, 2021), we strategically place the virtual tokens not at the very beginning of the source input but specifically before the target sequence, resulting in a final sequence structure of [source; prompt; target] (as described in Section 2.2.2). The application of soft prompting yields a 0.7% accuracy increase in the dense MediSwift-XL model, with significant gains also seen in the 50% and 75% sparse configurations. This highlights soft prompting’s role in refining outputs through task-specific conditioning, boosting accuracy in biomedical question answering tasks.
| MediSwift-XL | Size | Dense Fine-tune | Soft Prompt | Acc. |
| Dense | 1.21B | ✓ | ✗ | 76.1 |
| Dense | 1.21B | ✓ | ✓ | 76.8 |
| 605M | ✓ | ✗ | 75.7 | |
| 605M | ✓ | ✓ | 76.3 | |
| 302M | ✓ | ✗ | 74.8 | |
| 302M | ✓ | ✓ | 75.4 |
@@ 4.2 Document Classification on HoC
The performance is evaluated using the micro-F1 score, allowing direct comparison with prior models and demonstrating our method’s effectiveness. In Table 5, the dense MediSwift-XL model outperformed all similarly sized models in micro-F1 score, with its 50% and 75% sparse showing very competitive results, emphasizing sparse pre-training’s balance of computational efficiency and accuracy. This further showcases the potential of sparsity in optimizing language model performance for biomedical applications.
@@ 5 Related Work
Sparse weight training for language models (LM) have emerged as a promising avenue to address the computational intensity of training large models. Recent work has explored various sparse training methods (Thangarasa et al., 2023a; Dao et al., 2022a; Chen et al., 2022), aiming to maintain or enhance model performance while significantly reducing computational requirements. Techniques such as pruning (Chen et al., 2020), sparse activations (Mirzadeh et al., 2024), along with the development of specialized software (NeuralMagic, 2021; Gupta, 2024) and hardware (Lie, 2022b, a; Dietrich et al., 2021) have been pivotal. We build on these foundations, focusing on optimizing sparse weight training strategies specifically for domain-specific LMs, pushing the boundaries of efficiency.
@@
The evolution of language models for medical applications has progressed from adapting encoder-only architectures like BERT (Devlin et al., 2018), using biomedical corpora (Lee et al., 2019; Gu et al., 2021), to incorporating strategies like document links (Yasunaga et al., 2022b) and knowledge graphs (Yasunaga et al., 2022a). The shift towards autoregressive generative models, such as GPT (Brown et al., 2020b) and Llama (Touvron et al., 2023a), for pretraining on medical texts has led to significant advancements (Wu et al., 2023a; Luo et al., 2022; Bolton et al., 2021). Recent scaling efforts include GatorTronGPT with 20B parameters (Yang et al., 2022), as well as Clinical-Camel (Toma et al., 2023), MEDITRON (Chen et al., 2023) and Med-42 (Christophe et al., 2023), based on Llama-2-70B (Touvron et al., 2023b), focusing on mixed clinical and general English texts. Our work diverges from works that scaled up medical LMs by introducing weight sparsity into the pre-training of biomedical LMs. This reduces the computational costs typically associated with large-scale models, thereby improving the balance between efficiency and accuracy in the medical domain.
@@ 6 Conclusion
Limitations
Our work on MediSwift represents a significant leap forward in developing efficient domain-specific LLMs, particularly in biomedicine, by utilizing sparse pre-training to strike a fine balance between computational efficiency and accuracy. While we have initially focused on static sparse pre-training, the emerging field of dynamic sparse training (DST) holds great promise for further improvements (Evci et al., 2020; Mocanu et al., 2018; Liu et al., 2021a).
@@ Appendix B Unstructured Sparsity on Specialized Hardware Accelerators
Appendix C Author Contributions
We provide a summary of each author’s contributions:
Vithursan Thangarasa led the efforts on developing the method and pre-training the various dense and sparse MediSwift models. He also implemented several parts of the code for fine-tuning and proposed innovative approaches such as dense fine-tuning and soft prompting. Additionally, he wrote the manuscript, ensuring a comprehensive and cohesive presentation of the research.
Mahmoud Salem assisted in running experiments for both single-phase and multi-phase fine-tuning on PubMedQA. His work was crucial in validating the effectiveness of the proposed methods across different fine-tuning strategies.
Shreyas Saxena collaborated closely with Vithursan to help design the main method. His insights and feedback were vital in refining the overall approach and ensuring the robustness of the proposed techniques.
Kevin Leong contributed to the pre-training of both dense and sparse MediSwift models on the Cerebras CS-2. His expertise in utilizing the Cerebras CS-2 hardware was essential for the efficient and effective training of the models.
Joel Hestness was involved in discussions with Vithursan and provided valuable feedback to improve the manuscript. His contributions helped enhance the clarity of the manuscript.
Sean Lie frequently met with Vithursan for technical discussions and provided support for bringing up the Cerebras CS-2. His involvement was critical in performing the pre-training of MediSwift models from scratch and ensuring the smooth operation of the hardware.